### v1

# Shared Hardships Strengthen Bonds: Negative Shocks, Embeddedness and Employee Retention

Abstract
Unexpected events—“shocks”—are the motive force in explaining changes in embeddedness and retention within the unfolding model of labor turnover. Substantial research effort has examined strategies for insulating valued employees from adverse shocks. However, this paper provides empirical evidence that unambiguously negative shocks can increase employee retention when underlying firm and employee incentives with respect to these shocks are aligned. Using survival analysis on a unique data set of 466,236 communication records and 45,873 employment spells from 21 trucking companies, we show how equipment-related shocks tend to increase the duration of employment. Equipment shocks also generate paradoxically positive sentiments that demonstrate an increase in employees’ affective commitment to the firm. Our results highlight the important moderating role aligned incentives have in how shocks ultimately translate into retention. Shared hardships strengthen bonds in employment as in other areas.
Keywords: Worker Retention, Trucking, Turnover, Job Embeddedness
1 Introduction
Driver turnover for trucking companies routinely exceeds 90% per year (Goodman & Etheredge, 2022), costing trucking companies from \$7,894 (Suzuki, 2007; Trick et al., 2021) to \$15,705 per driver (Rodriguez et al., 1998) in 2024 dollars. Driver turnover increases freight prices in supply chains (Miller et al., 2020), with increases in transport costs ultimately borne by consumers (Carrière-Swallow et al., 2023). Higher turnover is also associated with higher accident rates (Miller et al., 2017; Shaw et al., 2005; Corsi & Fanara, 1988). Given the substantial direct and indirect costs, turnover is a perennial concern within the trucking industry(ATRI, 2023), and a continuing area of research in supply chain management and logistics.
In comparison to the national average voluntary separation rate of 2.2%, trucking industry turnover is extreme. [^1] December 2023; Latest figures: https://www.bls.gov/news.release/pdf/jolts.pdf These high rates of turnover provide variation that makes the trucking industry a particularly useful testbed for understanding successful worker retention strategies more generally. There are two important features of trucking, in particular, that make worker retention challenging. First, the job is highly-routinized across firms. Employees develop and depend little on firm-specific human capital so that their skills are highly substitutable between firms. In this environment, employers have little ability to offer firms-specific incentives to retain workers (Kryscynski et al., 2021). Second, the job is carried out remotely. While remote work, or work-from-anywhere policies have become important post-Covid, providing valuable amenities to workers and productivity gains to firms (Choudhury et al., 2021), this trend has a notable downside: it reduces employee embeddedness within firms because it makes it more difficult to build meaningful relational links with co-workers(Yang et al., 2022). These features constrain the strategies managers within the industry have for retention.
With truck drivers less tied to their job through relational linkages, they are more sensitive to job “shocks”–random and often unexpected events that cause employees to re-examine their current employment relationship (Lee & Mitchell, 1994; Mitchell et al., 2001; Lee et al., 2004). Additionally, deprived of relational options for embedding employees within an organizations, managers may be tempted to treat or prevent the shock, rather than addressing the underlying job-fit characteristics, which the shock highlights. For example, recent studies have found shocks such as pay (Conroy et al., 2022) or schedule(Bergman et al., 2023) variance increase turnover. A naive policy would be to defend against these shocks, such as by reducing pay frequency so that workers experience fewer of these negative events. However, while this is sometimes the only option for managers, it is the equivalent of palliative care, failing to treat the fundamental job-fit mis-alignments the shock highlights.
The main contribution of our paper is providing important empirical evidence that unambiguously deleterious events from either the firm or employee’s perspective, can result in increased retention and increased organizational commitment. We use data collected by a third-party human resources consultancy of weekly employee interviews of truck drivers for 21 trucking companies. We pair these weekly telephone interviews with employment records to follow over 45,873 employment spells. We aggregate communications records to identify 14 classes of shocks commonly faced by truck drivers, from equipment failures, and pay complaints, to issues of too little time-off. Using hazard analysis, we show that shocks where underlying incentives are aligned decrease driver turnover. In our sentiment analysis, equipment failures are followed by increased affective organizational commitment by drivers (Meyer & Allen, 1991). Our analyses use fixed effects and flexible hazard functions to control for heterogeneity across carriers, years, months, retention programs and agents. We also consider a number of alternative specifications and re-classifications of the shocks to demonstrate the robustness of our results.
To understand the interesting pattern, we borrow the unfolding model of labor turnover (Lee & Mitchell, 1994), where “shocks” cause employees to re-examine their current employment relationships. From the perspective of the employer, these shocks can be “positive” (reducing turnover), or “negative” (increasing turnover). Scare quotes are necessary because the ultimate effect of shocks is moderated by the extent to which an employee is embedded within a firm, that is, how the shock itself highlights existing job fit, relational linkages, and sacrifices (Lee & Mitchell, 1994; Mitchell et al., 2001; Lee et al., 2004). Harmful events with respect to other firm or employee objectives can increase retention. For example, an equipment failure is bad news, both for the firm and the driver. Drivers are paid piece-rate, so that inoperative equipment reduces their ability to earn. The firm, as residual claimant in this context, has similar incentives towards maintaining equipment at a high-level. When equipment fails, it highlights an important area of mutual interest alignment between the driver and the motor carrier organization, paradoxically increasing retention.
Our analysis yields a number of new and practical results. First, our extensive communications records allow us to assess the importance of a comprehensive set of work-related shocks on worker retention simultaneously. By far the most common, and the most strongly related to retention are equipment-related shocks (32%), and pay-related shocks (18%). Whereas shocks where interests are aligned, such as those related to equipment, tend to reduce turnover, pay-related shocks result in increased turnover precisely because interests are not aligned. We also contribute by performing a sentiment analysis allowing us to view the shocks from two distinct theoretical lenses (Philip & Medina-Craven, 2022): embeddedness theory (Mitchell et al., 2001), and organizational commitment(Meyer & Allen, 1991). Shocks that highlight interest alignment have positive associated sentiments, demonstrating increased commitment.
The remainder of the paper is organized as follows. Section 2 reviews selected job turnover research. Formal hypotheses are developed in section 3. Section 4 discusses the dataset and empirical model. Section 5 presents results. Section 6 concludes the paper, summarizing our results, discussing research implications, and suggesting areas for future research.
2 Literature Review
Worker retention and turnover have been studied extensively in management(Holtom et al., 2008), sociology (Granovetter, 2017), psychology (Hom et al., 2017) and economics (Rogerson et al., 2005). Worker turnover is costly, incurring not only direct costs in worker replacement, but can also be accompanied by productivity losses (Ton & Huckman, 2008), reduced performance (Li et al., 2022; Hancock et al., 2013), product reliability concerns (Moon et al., 2022) and knock-on effects through contagion (Porter & Rigby, 2021). Understanding and reducing voluntary employee separation (hereafter, “job turnover”), is therefore a continuing concern. Our research contributes to two streams of research: employee turnover (Hom et al., 2017) and, more specifically, to understanding turnover within the trucking industry (Conroy et al., 2022).
2.1 Job Turnover
Several meta-analyses synthesize turnover predictors, including (Park & Shaw, 2013; Zimmerman & Darnold, 2009). Predictors can be associated with job, training, managerial, individual, and macroeconomic characteristics. Well studied job-related turnover determinants include role clarity, job routinization, promotion opportunities, performance and pay. The job characteristics, when combined with managerial aspects (satisfaction with leadership, influence in decision-making, team cohesion) and work-environment characteristics (stress, communication, distributive justice), determine key employee attitudes, which are strongly predictive of voluntary turnover, including overall job satisfaction, and organizational commitment (Allen et al., 2010).
Nonetheless, there are a number of empirical challenges that have been difficult to overcome. Much of the turnover literature relies on exit surveys, which must then deal with issues of recall(Hom, 2011) and selection bias (Pecker & Fine, 2015), and has been shown to be less useful than attitude survey of current employees (Fottler et al., 1995). In particular, reports from former employees can be biased by post-decision rationalizations or overstated or made-up justifications (Griffeth & Hom, 2001; Westaby, 2005). Moreover, causal empirical assessment requires defining a counterfactual control group (Morgan & Winship, 2007), which is impossible in exit surveys where all employees are in the ”treated” group of exiting employees.
Our dataset, based on contemporaneous communications records from weekly-periodic driver checks-ins offers a number of advantages. First, the comparatively high frequency records reduce concerns of recall bias. They also allow us to examine heterogeneous response to shocks over time that hindsight misses. The evolving response to shocks is also theoretically important, particularly within the framework of the unfolding model of labor turnover (Lee & Mitchell, 1994; Lee et al., 1996), in which shocks are posited to have dynamic effects. Second, in contrast to studies relying on objective administrative records, our communications records add the color of employee perceptions and allow us to employ sentiment analysis to understand employee’s affective response to shocks.
Our paper is closely related with other empirical studies examining shocks and turnover such as Conroy et al. (2022); Bergman et al. (2023); Donnelly & Quirin (2006); Holtom et al. (2005). However, rather than focusing on specific shocks (Conroy et al., 2022; Bergman et al., 2023), we take and integrative approach examining a broad range of shocks, including pay (Conroy et al., 2022), scheduling (Bergman et al., 2023), time off (Westman & Etzion, 2001), training (Memon et al., 2016), health (Josephson et al., 2008), safety (Huang et al., 2016), and collegiality (Glambek et al., 2014), as well as others. The regularity and variety of shocks illustrates two important points for human resource managers seeking to limit turnover(Holtom et al., 2005): first, it is impossible to insulate employees completely against shocks; second, it is not always desirable to do so, as we show that harmful shocks can unexpectedly increase retention.
2.2 Trucking
Turnover is a particular concern in the supply chain, especially in trucking, where extreme turnover rates generate extensive industry  [^2] https://www.trucking.org/news-insights/truth-about-trucking-turnover and academic interest (Miller et al., 2021).
Studies have examined trucking turnover determinants that are common to other industries, for example, worker demographics, education and union status (Beilock & Capelle, 1990; Shaw et al., 1998), pay (LeMay et al., 1993; Phares & Balthrop, 2022; Conroy et al., 2022), tenure (Burks et al., 2006), economic climate (Prockl et al., 2017; Burks & Monaco, 2019; Miller et al., 2020), advancement opportunity (Min & Lambert, 2002), and fairness (Cantor et al., 2011). The logistics literature has also identified factors more particular to trucking, such as home time (Keller, 2002; Johnson et al., 2009, 2011; Burks et al., 2023) attitude towards dispatcher (Richard et al., 1995; Keller & Ozment, 1999; Morrow et al., 2005), lane balance (Sersland & Nataraajan, 2015) safety climate (Williams et al., 2011; Swartz et al., 2017) and emotional stress and loneliness (Kemp et al., 2013; Williams et al., 2017).
The self-stated purpose of much of this literature is to offer managerial insight, yet identifying antecedents to turnover is only the first step in designing programs to reduce turnover. Our study is the first to examine an explicit retention program, and is able to leverage a comparatively rich dataset on drivers’ self-reported concerns to uncover paradoxical correlations that highlighting the importance of interest alignment and job-fit in trucking.
3 Theory and Hypotheses
We use the unfolding model of labor turnover (Lee et al., 2004, 1996) as our primary theoretical scaffolding for hypothesis development. In this framework, labor turnover proceeds along four possible pathways, which for density of exposition, we present out of numeric order. It can occur because of evolving job dissatisfaction by the worker (path 4) or be precipitated by shocks–unexpected events that cause employees to re-evaluate their current employment relationships (paths 1-3). Path 2 is a “push” driven path without a pre-planned response and without a specific job alternative in mind. An example might be being assigned to a new sales region when a worker hasn’t had this experience previously. The shock triggers a re-evaluation of current employment relationship which can result in exit. Path 1 is a “script-driven” pathway, where an event occurs that triggers a planned and immediate exit. Examples include surpassing a savings target or being assigned to a new sales territory (when a worker has prior experience with this event) could trigger a worker to leave employment. Path 3 encompasses “pull” shocks where a specific job alternative enters consideration, as when an employee becomes aware of another job possibility. For our purposes, the usefulness of the unfolding model is not in making precise distinction between shock pathways 1-3, but in understanding what different types of exogenous shocks reveal about retaining employees.
Human resource managers can focus on two basic strategies for retaining employees: preventing or responding to shocks, or hardening employees against negative shocks. The latter strategy–the reason employees stay–is the complement of turnover, job embeddedness (Mitchell et al., 2001; Lee et al., 2014). Job embeddedness can be internal to the firm, or external, as when related to family, community, or other considerations. Dimensions of embeddedness applicable both internally and externally include links, fit and sacrifice.
As a class, truck drivers are much less embedded within a job than workers in other occupations because external embeddedness considerations rarely fix workers to a particular job. For example, a truck driver does not necessarily need to move locations to take a job with a different company, even if that company is located far away. Truckers are also less internally embedded. Drivers work remotely, making them less beholden to relational links that are important in other occupations. Additionally, skills are highly transferable across firms so that there is comparatively little sacrifice involved in changing jobs. Truck driver embeddedness relies more narrowly on considerations of job fit, resulting in more sensitivity to shocks.
Truckers experience a variety of shocks in the course of their duties which may trigger reassessment of current employment including traffic congestion, equipment failures, detention during loading and unloading, variation in pay, and so on. The effect of the shock, however is moderated by embeddedness, and, in particular, what the shock reveals about job-fit. The ultimate effect of a shock, with respect to retention, is predictable based on interest alignment between firm and employee.
There are shocks in trucking where drivers and employer interests are aligned. Equipment maintenance is one example. Long-haul truck drivers live out of their vehicles so that keeping vehicles in good condition contributes substantially to driver welfare. Drivers are liable for any maintenance problems that result in potential safety hazards. Additionally, because drivers are often paid piece-rate by mileage, any equipment failures that prevent vehicle operation cuts into the drivers’ pay. Drivers therefore have the incentive to immediately address and report equipment-related problems.
Correspondingly, employers have the incentive to immediately respond and take substantive action to fix equipment-related problems. First, equipment-related problems that prevent a vehicle from operating hurt firms through reduced operational efficiency and potential late-delivery penalties from shippers. Second, severe equipment problems may result in safety violations, negatively affecting a firm’s safety record and thus their ability to do business.
We hypothesize, then, that equipment-related shocks underscore employee-employer interest alignment, which reduces turnover. Stated formally, we then have H1:
Hypothesis 1.
Shocks which highlight job-fit through interest alignment increase embeddedness, therefore reducing turnover.
Other examples of shocks where interests are aligned are in resolving operational inefficiencies, such a driver detention, inefficient routing, and so on.
Driver and employer interests are not aligned for all shocks. Pay and home time provide important counter-examples (Burks et al., 2023). Piece-rate pay per mile implies that any pay increase for drivers reduces the residual for the firm; moreover, pay increases have not been shown to increase productivity in this industry (see, for example,(Belzer & Sedo, 2018) on income targeting). More home-time for drivers, means less capacity for the firm and lowered capital utilization when drivers take vehicles home (Burks et al., 2023), Conversations between employers and employees about pay and home time are thus more likely to be acrimonious and less likely to result in fruitful resolution. The resultant interest divergence reduces embeddedness, increasing turnover. Formally, we have H2:
Hypothesis 2.
Shocks which expose interest-divergence increase turnover.
4 Data and Empirical Model
4.1 Data
Our empirical analysis uses communication records from a human resource outsourcing company specializing in retention programs for trucking companies. The core of their programs is making weekly (or longer interval) phone calls to truck drivers to identify and record their “issues”– “shocks” in the terminology of the unfolding model of turnover– and then pass them to their customers (i.e., trucking companies) to help them to be proactive in driver retention. Specifically, the retention programs typically last for three to six months for each driver and during the program, the agents of the company record a brief summary of their conversation after each call and tag it with proprietary classification codes for the issues that they identify during the conversation (hereafter issue codes). Some examples of the communication notes are below:
  • Good afternoon Mr. …. All is well sir the miles are rolling in well the road has been good thank you. Thank you for checking in. Be well!
  • Not great. She decided to quit as she was messed over on her apprenticeship hours and all that. She hated that they sent her all the way to NY for her training.
The classification system includes more than two hundreds issue codes; we re-classify these codes into fourteen major categories: Home-time, Pay, Equipment, Safety/Ethics, Training, Team, Operations, Disrespect, Company, Covid, Waiting, Personal, Health, Others. [^3] Due to an non-disclosure agreement with the human resource company, the full list of issue codes and their definitions can’t be shared here. Instead, we include some examples for each issue category in Appendix A. In our main analysis, we further aggregate the categories into two: “Aligned,” where employee-employer interests are aligned, including Equipment, Operations and Waiting, and “Misaligned,” where their interests are opposed, which includes all other categories.
We merge these communication records with driver employment data to create survival-time data suitable for duration analysis. The employment data includes hire and termination dates, from which we calculate employment duration (in weeks) and termination status at the time of each communication record. For drivers whose employment was not terminated during the program, no termination date is observed and their employment spells are treated as censored (i.e., right-censoring). To create a more homogeneous sample, we exclude non-voluntary termination cases (i.e., disqualified quit) and employment spells longer than 100 weeks. [^4] Our results in the next sections are robust to this sample selection. The data is then merged with employer data such as carrier-level scores related to unsafe driving, hours-of-service compliance, vehicle maintenance, misuse of controlled Substances/alcohol from Safety Measurement System (SMS) of the Federal Motor Carrier Safety Administration.
The duration and termination status variables, along with the communication records, are the key variables in our duration analysis. Our data is multi-record-per-subject survival data, where individual drivers are observed multiple times with different communication records over their employment spell. This provides our analysis additional strength because it allows for time-varying regressors that can account for the entire history of covariates. This feature is particularly useful for analyzing the effects of lagged variables.
Our final sample includes 466,236 communication records (= total observations) of 45,873 employment spells from 21 carriers over 2019 - 2022. Most spells are from large carriers, with the largest carrier accounting for 60% and other seven carriers accounting for 32.6% of the sample. [^5] Employment concentration at larger firms is typical in this industry. While the majority of firms are small, the power-law distribution of firms size dictates that the majority of workers work for larger firms (Balthrop, 2021) Later years includes more observations: 16,882 (2019), 71,839 (2020), 158,221 (2021), 219,294 (2022). [^6] These are the numbers of communication records from the spells initiated in a given year. The sample includes four different retention programs, four call-intervals, and fifteen agents, where two programs account for 97.8% of the sample and most observations are based on weekly or bi-weekly phone calls (99.8%). As we shall see later, our duration analysis controls for heterogeneity across carriers, years, months, programs, call-intervals and agents using fixed effects and flexible hazard functions. [^7] In practice, the retention programs and call-intervals are determined by carriers, and in the data, program type and call-interval do not vary much within carriers (i.e., each carrier applies the same program and call-interval to all their drivers). This indicates that controlling for carrier-level heterogeneity is particularly important and an effective way to control for overall heterogeneity in our data.
The distribution of the issue codes under the fourteen major categories is provided in Table 1 below. The most frequent issue was Equipment, followed by Pay, Disrespect and Safety/Ethics. In terms of the “Aligned” (i.e., Equipment, Waiting and Operations) and “Misaligned” (i.e., the rest) categories, about 47% of the codes issued belong to the “Aligned” category.
[$=$ Table 1 here $=$]
Next, we provide Kaplan-Meier (KM) survival curves in Figure 1 to summarize some important features in the duration data. The KM survival curves non-parametrically estimate the survival rates at each employment week (i.e., the proportion of employment spells that were not terminated after a given number of weeks). The shaded areas represent 95% confidence intervals. We estimate the curves by carrier and year to highlight some of important heterogeneity in the data. The carriers are ranked in terms of the number of employment spells and their rankings are denoted by “Top-.”
[$=$ Figure 1 here $=$]
The survival curves demonstrate the high turnover rate in trucking industry. Top1 and Top2-8 carriers’ survival rates are below 0.5 after 24 weeks, indicating that more than 50% of the drivers in these carriers quit their job within six months. A significant heterogeneity across carriers is observed. The survival rate of Top2 carrier is significantly higher than the other carriers, with the gap widening as the employment duration increases. The heterogeneity across years is also interesting. The survival curve of year 2019 flattens around fifteenth week, implying that driver turnover significantly decreases after the period. However, after 2020, turnover worses, even for long-duration employees.
4.2 Empirical Model: Cox Proportional Hazard Model
We use the Cox Proportional Hazard (PH) model to analyze the effects of issue types on driver turnover. The Cox PH model specifies the hazard function, i.e., the probability that an employment spell is terminated at $t$ given that it has not been terminated before $t$, [^8] More precisely, in case of discrete time, the hazard function without covariates is defined as $\lambda(t)=\Pr[T=t|T\geq t]$ (= conditional failure probability), where $T$ denotes the duration random variable, whereas in case of continuous time, it is defined as $\lambda(t)=\lim_{h\to 0}\Pr[t\leq T\leq t+h|T\geq t]/h$ (= instantaneous failure rate), which is technically not a probability, but in practice, it is loosely interpreted as “instantaneous” probability. as
$\lambda(t|\mathbf{x})=\lambda_{0}(t)\exp\left(\mathbf{x}^{{}^{\prime}}\beta\right),$ (1)
where $\lambda_{0}(t)$, called the baseline hazard, is a function of $t$ (duration) alone and $\exp\left(\mathbf{x}^{{}^{\prime}}\beta\right)$ is a function of $\mathbf{x}$ (a vector of covariates) alone, so the two functions separate the hazard rate into two portions that varies by $t$ and that varies by $\mathbf{x}$. The separability is the key assumption in the Cox PH model and allows the estimation of $\beta$ without specifying the baseline hazard in the partial likelihood estimation approach proposed by Cox (1972, 1975). The baseline hazard function can be recovered after estimating $\beta$ using a method similar to the KM estimator. In duration models, the hazard function, which fully characterizes the distribution of duration, is the main building block that models the associations between duration and covariates of interest.
[…]
We use fixed effects and flexible hazard functions to control for heterogeneity across carriers, years, months, programs, call-intervals and agents. In all specifications, we include fixed effects for year, month, retention program, call-interval and agent. Since much of the retention program’s details (length, call-interval, etc) are determined by carriers, and carrier characteristics are arguably the most important factors in driver’s quit decision, we employ three different approaches to control for heterogeneity across carriers: fixed effects, carrier-level variables from the SMS, and carrier-specific baseline hazard functions. In the result tables in the next sections, Model 1 and 2 includes carrier fixed effects and carrier-level variables, respectively, and Model 3 allows for carrier-specific baseline hazard functions. [^12] The partial likelihood estimation of the Cox PH model calculates the likelihood of each (observed) termination event by the hazard rate of the spell terminated at the event divided by the sum of the hazard rates of the terminated and non-terminated (i.e., censored) spells existed at the event. The division in the calculation cancels out the baseline hazard function due to the specification in (1), which allows us to estimate $\beta$ without specifying the baseline hazard function. When group-specific baseline hazard functions are allowed, the ”partial” likelihoods are calculated by group.
5 Results
5.1 Effects of Issue Types on Driver Turnover
To examine our hypotheses, Table 2 reports the coefficient estimates for “Aligned” and “Misaligned” issues in the Cox PH model. [^13] Remember that the estimate for each issue type is the average of the estimates for its original issue dummy and four lags, so it includes the dynamic effects. As defined in Section 4.1, “Aligned,” where employee-employer interests are aligned, includes Equipment, Operations and Waiting, and “Misaligned,” where their interests are opposed, includes all other categories. Coefficients estimates are stable across specifications.
[$=$ Table 2 here $=$]
A stark contrast between the two types of issues is observed: all other things being equal, “Aligned” issues on average decrease driver turnover, whereas “Misaligned” issues increase it. “h.r.” are hazard ratios, which helps the reader quickly appreciate the magnitude of the impact. For example, in Model 1, the hazard ratio of Aligned is $-0.29$, indicates that experiencing a shock that highlights aligned interests decreases the hazard (i.e., instantaneous probability of leaving the job) by 25% over the baseline hazard. On the contrary, the hazard ratio for Misaligned issues is $0.32$, indicates having an issue like Home-time, Pay, etc increases the hazard by nearly 38%. The somewhat counter-intuitive result here is that Aligned issues lower driver turnover, even though they are still negative events that drivers experienced. For example, a driver that experiences problems with having to wait at a loading dock is more likely to remain at the motor carrier than a driver that experiences no wait-time shock.
We argue that this is because the Aligned issues provided opportunities for both drivers and employers to communicate with positive messages and confirm their mutual interests, thus increasing the employee’s embeddedness. When the driver experiences operational difficulty, equipment failure, or extended wait times, these shocks also harm the firm by directly reducing worker and equipment utilization. A worker detained at a loading dock, or broken down on the roadside, is not transporting freight. In these cases the motor carrier as an organization has incentive to sympathize with workers, which may result in strengthened relational linkages between frontline workers (drivers) and their support teams. Beyond words, the firm is also likely to take action to remedy the problem. Whether through words (which we observe) or actions (unobserved) these Aligned shocks highlight the firm and workers’ common purpose, while potentially revealing the firm’s support infrastructure available to the worker.
We further disaggregate these results in Table 3, which reports the coefficient estimates for fifteen major issue types in the Cox PH model. [^14] More detailed results including the estimates of individual lag terms are available upon request.
[$=$ Table 3 here $=$]
As with our main results, coefficient estimates in Table 3 are not sensitive to specification, whether carrier fixed effects (Model 1), carrier-level variables (Model 2) or carrier-specific baseline hazards (Model 3) are included. Categories most strongly associated with increased turnover include Covid and Personal issues. While Covid is a strong indicator of potential exit, it is a comparatively rare issue (only 224 instances), even though our sample period covers the onset of the pandemic. Personal issues are more common, and have stronger effects on turnover. This category encompasses issues related to family, which supports previous research establishing loneliness and family disruption as one of the most important reasons for truck driver turnover (Williams et al., 2017). Even the smaller (but positive) coefficients have managerially meaningful effects on hazard: ethics complaints increase the baseline hazard rate by 25% (model 1). In contrast, the aligned issues, including Equipment, Operations, and Waiting, all result in strongly reduced hazard rates.
A potential concern, however, is that there may be unobserved driver types with different hazard rates and these types are correlated with the type of issues they reported. For example, there may be “attached” and “detached” drivers with low and high turnover rates, and the former report aligned issues more frequently than the latter, simply because they are more attached to their company. Then, the signs of the coefficients on Aligned and Misaligned issues are simply a reflection of the hazard rate difference between the two types, not the effects of the two different types of issues. Said differently, it is possible to imagine a driver experiencing an equipment issue but not reporting it precisely because the driver expects the issue will go unresolved. Such a disillusioned driver can be expected to have a higher turnover rate; importantly, this type of detached driver will fall in the base-case “no issue” category, biasing base-case turnover rates upward.
To address this potential endogeneity issue, we employ two approaches. First, we create a variable quantifying the responsiveness of each driver to the human resource company’s calls. Responsiveness is an important revealed characteristic that can be correlated with driver types. We calculate the proportion of the calls for each driver tagged with “voice mail,” “no answer,” etc and include it into the duration models. The call answer rate is quite uniformly distributed across drivers, with quartiles 0.20, 0.42, and 0.64. The results are reported in Table B1 in Appendix B, where the estimates largely remain the same.
Next, we focus on the sub-population of communicators. That is, rather than making comparison to a reference group that may under-report the shock, we make comparisons between groups that report multiple shocks. First, we dis-aggregate the Aligned category into “Broken-equipment” and “Other aligned,” where “Broken-equipment” includes issues related to “broken/missing components” only, which is arguably less subject to the under-reporting problem. Second, we re-classify the issues into new categories that account for joint-distribution of Aligned and Misaligned issues: Broken-equipment only (B-only), Other aligned only (OA-only), Misaligned only (MA-only) and their combinations. [^15] Note that a communication record can be tagged with multiple codes in difference categories. Communication records counted multiple times for the fourteen categories in Table 1 can be counted fewer times for the new categories in Table 4, resulting in different total numbers in the tables.
We note that, if other covariates are ignored, the coefficient on “Aligned” in Table 2 is calculated by a combination of two hazard rate differences: a hazard rate difference between the case experiencing Aligned issues only and the case experiencing no issues, and another hazard rate difference between the case experiencing both Aligned and Misaligned issues and the case experiencing Misaligned issues only. Roughly speaking, the first difference compares drivers reporting Aligned issues only and drivers who were silent, whereas the second difference compares drivers who report at least one of the two types of issues. Because the second comparison is among “communicators,” it is arguably less subject to the under-reporting issue. The new categories allow us to identify these dis-aggregated effects.
[$=$ Table 4 here $=$]
Table 4 include the distribution of the issues under the new categories, and Table 5 reports the coefficient estimates for the new categories. The coefficient on “Broken & Misaligned” subtracted by the coefficient on “Misaligned” calculates the effect of Aligned issues that is less subject to the under-reporting issue as argued above, which is -0.14 in Model 1 (i.e., specification including carrier fixed effects), with standard error 0.08, t-statistic -1.78 and p-value 0.07.161616The standard error is calculated using the delta method. The results from Model 2 and 3 specifications are qualitatively the same. The p-value is above 5%, but the result overall indicates that relative to the population experiencing only Mis-aligned shocks, a driver experiencing both an Aligned and Mis-aligned shock has a lower hazard rate. The striking result that two negative shocks are better than one shock provides strong support for our hypotheses, at least in this sub-population.
[$=$ Table 5 here $=$]
5.2 Potential Mechanism: Sentiment Change after Issues
To explain the potential mechanisms for the turnover rate changes observed in the previous section, we employ sentiment analysis to examine drivers’ affective response. We use the VADER (Valence Aware Dictionary and sEntiment Reasoner) algorithm to compute sentiment scores, ranging -1 to 1 (the larger, more positive), for each communication note in the data. The VADER algorithm is a popular lexicon and rule-based sentiment analysis tool, which combines a sentiment lexicon–a list of words labeled as either positive or negative according to their semantic orientation–with certain rules (e.g., how to handle negation, modifier, etc) to compute sentiment scores. Compared to other sentiment analysis tools (e.g.,TextBlob), the VADER is specifically attuned to sentiments expressed in microblog-like contexts (e.g., social media) (Hutto, 2014), which is appropriate for our analysis given the form and style of the communication notes in our data. [^17] For sentiment scores, we particularly use the compound score in the VADER, which combines positive, neutral and negative scores calculated in the algorithm.
We regress the sentiment scores on the same set of regressors used in the duration models in Table 5, and plot the coefficient estimates and their confidence intervals, including the lag terms, in Figure 2.181818The full set of results is contained in Table C1, Appendix C. Note that the scale of y-axis in the second panel of Figure 2 is adjusted to show the changes in coefficients more clearly. Driver issues all have a negative effect on the sentiment of conversations right after the issues. Because it is a common practice to classify sentences as positive, neutral or negative if the VADER sentiment score is greater than 0.05, between 0.05 and -0,05, or less than -0.05, respectively,191919 https://github.com/cjhutto/vaderSentiment#about-the-scoring the sizes of the coefficients at period = 0 (i.e., the period issues were identified) implies a substantial negative effect of driver issues on conversation sentiment.
However, the pattern of how the negative sentiment changes over time is significantly different across issue types. In case of “Broken-equipment only,” the negative sentiment turns to positive after one period and the positiveness continues. On the other hand, the negative sentiment on “Misaligned only” remains negative. The coefficients on the combinations of the three issue types are less significant due to smaller numbers of such cases, but a similar pattern is observed in the second panel of Figure 2: as the case is more associated with “Aligned” issues, the negative effect tends to disappear faster. These observed relationships are consistent with companies taking remedial action in the event of broken equipment. It is plausible that the lagging positive sentiments are then the response to such actions.
6 Discussion
Employees experience numerous and varied shocks which cause them to reflect upon and re-evaluate their current employment relationship. This paper catalogs self-reported shocks from a sample of over 40,000 truckers, which we categorize into Aligned and Misaligned categories. The most commonly reported events are problems related to equipment (Aligned) followed by pay complaints (Misaligned). The surprising result is that Aligned shocks, other things equal, result in longer job spells.
This finding is understandable within the framework of the unfolding model of labor turnover (Lee & Mitchell, 1994). Shocks are not direct determinants of retention, instead they highlight specific aspects of the current employment relationship in terms of links, fit, and sacrifice. Aligned shocks have the ultimate effect of underscoring job-fit compatibility between employee and employer. Motor carriers have the incentive to address Aligned issues immediately, leaving the employee feeling listened to and supported, thereby increasing embeddedness and ultimately retention. Indeed, sentiment analysis demonstrates positive employee feelings directly following equipment related shocks. Mis-aligned shocks have the opposite effect. Rather than highlighting job fit through interest alignment, these complaints underscore fundamental interest divergence at the current job.
While pay is one of the most common complaints for truck drivers, its effect on retention is weak in comparison with home time, company and personal issues. While home time has been previously highlighted (Keller, 2002; Johnson et al., 2009, 2011), the company effects have not. This category includes policies with regards to pets, driver-facing cameras, and sleep apnea machines; however, the most common complaint in the category is the inability to take the tractor home at the end of a shift. Given that drivers’ homes, equipment yards and load locations can be separated by hundreds of miles, it is easy to understand the importance of this amenity. Managers may be willing to accept lower equipment utilization rates in favor of increased retention. Within the “personal” category the most important issue for drivers is being away from family. Offering more amenities in terms of family support is another important measure managers can focus on to increase retention.
The sheer variety of shock makes it impossible to entirely insulate employees. But negative shocks can have a silver lining when they underscore interest alignment. Our research uncovers a potentially under-used strategy for managers, which is to allow negative shocks to happen, but to address them immediately. In our context, the act of rescuing an employee from an equipment failure increases retention. Shared hardships strengthen the bonds of committment.
Appendices
Appendix A Fifteen Major Issue Categories and their Examples
  • Home-time: broken promises, shift change
  • Waiting: load/unload delays, waiting on assigned truck
  • Pay: inaccurate, not competitive
  • Equipment: broken/missing components, maintenance issues
  • Safety/Ethics: improperly loaded, unsafe operations
  • Training: too many hours, waiting on assigned truck
  • Team: co-driver issues, waiting on co-driver
  • Operations: fueling issues, unloading issues
  • Disrespect: slow/no response
  • Company: sleep apnea policy, pet policy
  • Covid: physically unable to work
  • Personal: family reasons
  • Health: health issues
  • Others: no show
Appendix B Robustness Check with Call Answer Rate
The same data and specifications in the models in Table 3 are used except for including an additional variable quantifying the responsiveness of each driver to the human resource company: the proportion of calls tagged with “voice mail,” “no answer,” etc. The log-pseudo-likelihood values of Model 1, 2, and 3 are -210,952.6, -211,002.9, and -184,545.8, respectively.
Table B1: Robustness Check with Call Answer Rate
Category Model 1 Model 2 Model 3 Category Model 1 Model 2 Model 3
Home-time 0.38 0.38 0.38 Operations -0.20 -0.20 -0.20
(s.e.) (0.03) (0.03) (0.03) (0.05) (0.05) (0.05)
[[[[t-stat]]]] [[[[10.89]]]] [[[[11.00]]]] [[[[10.92]]]] [[[[-4.25]]]] [[[[-4.22]]]] [[[[-4.26]]]]
{h.r.} {0.46} {0.46} {0.46} {-0.18} {-0.18} {-0.18}
Waiting -0.17 -0.16 -0.17 Disrespect 0.10 0.10 0.10
(0.04) (0.04) (0.04) (0.03) (0.03) (0.03)
[[[[-4.12]]]] [[[[-3.86]]]] [[[[-3.97]]]] [[[[2.98]]]] [[[[3.00]]]] [[[[2.97]]]]
{-0.16} {-0.15} {-0.15} {0.11} {0.11} {0.11}
Pay 0.22 0.22 0.22 Company 0.27 0.26 0.28
(0.02) (0.02) (0.02) (0.09) (0.09) (0.09)
[[[[9.65]]]] [[[[9.44]]]] [[[[9.71]]]] [[[[3.04]]]] [[[[2.94]]]] [[[[3.25]]]]
{0.25} {0.24} {0.25} {0.31} {0.30} {0.33}
Equipment -0.25 -0.24 -0.25 Covid 0.44 0.44 0.42
(0.02) (0.02) (0.02) (0.13) (0.13) (0.14)
[[[[-11.73]]]] [[[[-11.48]]]] [[[[-11.70]]]] [[[[3.26]]]] [[[[3.26]]]] [[[[3.12]]]]
{-0.22} {-0.22} {-0.22} {0.55} {0.55} {0.53}
Safety/Ethics 0.22 0.23 0.21 Personal 0.60 0.59 0.60
(0.05) (0.05) (0.05) (0.04) (0.04) (0.04)
[[[[4.67]]]] [[[[4.96]]]] [[[[4.52]]]] [[[[14.94]]]] [[[[14.74]]]] [[[[14.97]]]]
{0.25} {0.26} {0.24} {0.82} {0.81} {0.83}
Training 0.24 0.23 0.24 Health 0.32 0.32 0.31
(0.04) (0.04) (0.04) (0.06) (0.06) (0.06)
[[[[5.43]]]] [[[[5.13]]]] [[[[5.32]]]] [[[[5.26]]]] [[[[5.40]]]] [[[[5.01]]]]
{0.27} {0.26} {0.27} {0.37} {0.38} {0.36}
Team 0.08 0.07 0.09 Others 1.08 1.06 1.16
(0.09) (0.09) (0.09) (0.21) (0.19) (0.18)
[[[[0.93]]]] [[[[0.80]]]] [[[[1.00]]]] [[[[5.23]]]] [[[[5.47]]]] [[[[6.35]]]]
{0.08} {0.07} {0.09} {1.94} {1.89} {2.17}
Appendix C The Full Sentiment Regression Result
The outcome variable is the VADER (Valence Aware Dictionary and sEntiment Reasoner) score. The regression models include the same sets of covariates and fixed effects in the models in Table 5. Standard errors are clustered at the spell level.
Table C1: Change of Conversation Sentiment after Driver Issues
Category coeff. s.e. t-stat
Broken-equipment only
t=0 -0.051 0.006 -8.560
t=1 0.009 0.005 1.920
t=2 0.020 0.005 4.340
t=3 0.017 0.005 3.530
t=4 0.021 0.005 4.210
[…]
t=0 -0.057 0.006 -9.040
t=1 -0.002 0.005 -0.370
t=2 0.007 0.005 1.370
t=3 -0.007 0.005 -1.320
t=4 0.005 0.005 1.030
[…]
t=0 -0.049 0.004 -11.480
t=1 -0.017 0.003 -5.050
t=2 -0.015 0.003 -4.290
t=3 -0.009 0.004 -2.470
t=4 -0.005 0.004 -1.310
[…]
t=0 -0.233 0.013 -18.310
t=1 -0.023 0.009 -2.410
t=2 -0.010 0.009 -1.070
t=3 -0.021 0.010 -2.140
t=4 -0.005 0.010 -0.530
[…]
t=0 -0.237 0.022 -11.010
t=1 -0.025 0.016 -1.600
t=2 -0.024 0.016 -1.480
t=3 -0.004 0.017 -0.210
t=4 -0.011 0.018 -0.600
[…]
t=0 -0.214 0.011 -18.910
t=1 -0.043 0.008 -5.030
t=2 -0.026 0.008 -3.070
t=3 -0.024 0.009 -2.730
t=4 -0.019 0.010 -1.930
[…]
t=0 -0.381 0.031 -12.120
t=1 -0.076 0.023 -3.360
t=2 -0.051 0.024 -2.080
t=3 -0.047 0.024 -1.990
t=4 -0.015 0.025 -0.620